A particle is moving along a curved path. At a certain instant when the slope of the 

path is 0.75, ax = 6 fps2 and ay = 10 fps2

. Compute the values of the total acceleration

(a), tangential acceleration (at) and normal acceleration (an) at this instant.



1
Expert's answer
2021-05-05T07:40:50-0400

a)


"a=\sqrt{6^2+10^2}=11.7\frac{ft}{s^2}"

b)


"a_t=11.7(0.75)=8.8\frac{ft}{s^2}"

c)


"a_n=\sqrt{11.7^2-8.8^2}=7.7\frac{ft}{s^2}"